OpenStudy (anonymous):

triangle ABC IS circumscribed about circle O , and D, E , and F are points of tangency. AF=10,CE=20, and BD=30 , Find the lengths of the sides of the triangle and show that the triangle is right triangle.

OpenStudy (anonymous):



OpenStudy (mertsj):

You know the lengths of the sides. If it is a right triangle, the Pythagorean Theorem will prove that it is.

OpenStudy (anonymous):

apply Pythagoras Theorem to this triangle i.e. AB^2+AC^2=BC^2 and here 30^2+40^2=50^2 , 2500=2500. This shows that the sides make pythagorean triplet so this is a right triangle.